OpenStudy (anonymous):

Someone please help!!! find the zeros of the polynomial function and state the multiplicity of each f(x)=3(x+8)^2(x-8)^3

myininaya (myininaya):

find when f is 0.

myininaya (myininaya):

\[0=3(x+8)^2(x-8)^3 \]

myininaya (myininaya):

\[=>(x+8)^2=0 \text{ or } (x-8)^3=0\]

myininaya (myininaya):

The first one there is suppose to be two answers because of the square. The second on there is suppose to be three answers because of the cube.